# Engagement Maximization

### v5

@@ A.6 Proof of 5
We slightly abuse notation and let $x_{1},x_{2}$ denote the random variables that determines the state realization. $\forall\pi$, let $\pi_{i}=\pi|_{X_{i}}$. Let $s_{1}$ be the random variable that is independent to $x_{2}$ and induces posterior belief $\pi_{1}$ for $x_{1}$. Let $s_{2}$ be the random variable that induces posterior belief $\pi$. Let $I^{M}$ denote mutual information. Then,
@@
Therefore, let $\pi^{\prime}=\pi_{1}\otimes\pi_{2}$,
@@
$\mathbb{E}^{\pi}[\hat{u}(q)-\hat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi_{1}}[\hat{u}(q,\bar{q}_{0}|_{X_{2}})-\hat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi^{\prime}}[\hat{u}(q)-\hat{u}(\bar{q}_{0})];$
@@
$\mathbb{E}^{\pi_{1}\otimes\pi_{2}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)+\lambda G(q)\right]$
$=$ $\underbrace{\mathbb{E}^{\pi_{1}}\left[\hat{u}(q,\bar{q}_{0}|_{X_{2}})-\frac{\bar{\kappa}}{\chi}H(q,\bar{q}_{0}|_{X_{2}})\right]}_{\text{maximized by $\pi^{*}_{1}$ solving $V^{B}$}}+\mathbb{E}^{\pi_{2}}\left[-\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0}|_{X_{1}},q)+\lambda G(\bar{q}_{0}|_{X_{1}},q)\right].$
Appendix B Omitted Proofs for Section 4
B.1 Proof of 2
@@
Sufficiency: condition (ii) of Definition 3 is trivially satisfied. Condition (iii) is a direct implication of Propositions 1 and 2. Next, we verify condition (i). Suppose for the purpose of contradiction that condition (i) is violated, i.e. for some $t$ and $q\in\text{Supp}(q_{t}|t<\tau)$,
$\mathbb{E}^{P}[G(q_{\tau})-G(q)|t<\tau]<J(q);$
Let $(P^{\prime},I^{H}{}^{\prime},\tau^{\prime})$ be the policy that implements $J(q)$. Note that
$\mathbb{E}^{P^{\prime}}[G(q^{\prime}_{\tau^{\prime}})-G(q)|\mathcal{F}^{\prime}_{0}]>\mathbb{E}^{P}[G(q_{\tau})-G(q)|t<\tau];$
$\mathbb{E}^{P^{\prime}}[\hat{u}(q^{\prime}_{\tau^{\prime}})-\bar{\kappa}(\tau^{\prime}+t)|\mathcal{F}^{\prime}_{0}]\geq\hat{u}(q)-\bar{\kappa}t=\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|t<\tau].$
Therefore, replacing the policy $(P,I^{H},\tau)$ by $(P^{\prime},I^{H}{{}^{\prime}},\tau^{\prime}+t)$ from the positive probability history $(q_{t}=t,t<\tau)$ strictly improves the principal’s payoff and weakly improves the agent’s payoff, leading to a contradiction.
Necessity: Suppose for the purpose of contradiction that the equality doesn’t hold, then, there exists $t$ and $q\in\text{Supp}(q_{t}|t<\tau)$ s.t.
$\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|t<\tau]\geq\hat{u}(q)-\bar{\kappa}t>0.$
We prove that this violates condition (i), which states
$\mathbb{E}^{P}[G(q_{\tau})-G(q)|t<\tau]\geq J(q).$
However, this is impossible as the policy $q_{\tau}|t<\tau$ leaves the agent with positive surplus while Theorem 1 implies that the agent’s participation constraint is always binding for the policy implementing $J(q)$. ∎
Online Appendix to ‘Engagement Maximization’
Appendix C Omitted Proofs for Section 5
C.1 Proof of Theorem 3
Proof.
By Assumption 3, it is w.l.o.g. to consider strategies whose marginal distribution of $(q_{\tau},\tau)$, which we denote $F(q,t)\in\Delta(\Delta(X)\times\mathbb{R}_{+})$, is rotationally symmetric, because any implementable process $q_{t}^{*}$ can be rotated by $R$ and remains implementable. Therefore, $\mathbb{E}^{F(q,t)}[q|t\geq t^{\prime}]$
$=\mathbb{E}^{F(q,t)}\left[\sum_{j=0}^{J-1}\frac{1}{J}R^{j}q\big{|}t\geq t^{\prime}\right]=q_{0}$ for all $t^{\prime}\geq t_{0}$. Now, consider the following relaxed problem for the principal:
$\sup_{F\in\Delta(\Delta(X)\times\mathbb{R}^{+})}$ $\rho\int_{\mathcal{P}(X)}\int_{0}^{\infty}H(q)F(dq,dt)$ (7)
[…]
Let $F^{*}$ be the joint distribution of $(q_{\tau^{*}}^{*},\tau^{*})$ under the policies $(P^{*},I^{*},\tau^{*})$. We prove that $F^{*}$ solves equation 7, and hence that the policies $(P^{*},I^{*},\tau^{*})$ achieve weakly higher utility than any other policies.
We first write down the Lagrangian: let $\lambda(t)$ and $\gamma(t)$ be the multipliers of the first and second constraints, respectively,
[…]
$-\int_{\mathcal{P}(X)}\int_{0}^{\infty}\Lambda(t)H(q)F(dq,dt)$
$+\int_{\mathcal{P}(X)}\int_{0}^{\infty}\Gamma(t)\widehat{\mu}(q)F(dq,dt)-\int_{0}^{\infty}\left(\int_{0}^{t}\gamma(s)K_{s}(t)\right)F(dt)$
$+\int_{\mathcal{P}(X)}\int_{0}^{\infty}\xi(q,t)F(dq,dt)+\nu(1-\int_{\mathcal{P}(X)}\int_{0}^{\infty}F(dq,dt)),$
[…]
Note that the characterization applies to the constant $\kappa$ case as well: there exist $\Lambda_{t}$ and $\Gamma_{t}$ (both are constant fixing $t$) s.t.
$\begin{cases}\nabla H(q)(\rho-\Lambda(t))+\Gamma(t)\nabla\widehat{\mu}(q)=0\text{ $\forall q$ in the support of $\pi^{*}(\kappa(t))$}\\ \chi\Lambda_{t}-\Gamma_{t}\kappa(t)=0\end{cases}$
The first equality is from concavification. The second equality is from the FOC. Now define $\Lambda(t)$ and $\Gamma(t)$:
$\begin{cases}\frac{\rho-\Lambda(t)}{\Gamma(t)}=\frac{\rho-\Lambda_{t}}{\Gamma_{t}}\\ \chi\Lambda(t)-\Gamma(t)\kappa(t)+\gamma(t)\mathbb{E}^{\pi^{*}(\kappa(t))}[\widehat{\mu}(q)]+\lambda(t)E_{\pi^{*}(\kappa(t))}[H(q)]=0\end{cases}$
Replace $\Lambda(t)$ with $\rho+\left(\frac{\kappa(t)}{\chi}-\frac{\rho}{\Gamma_{t}}\right)\Gamma(t)$, we get an ODE for $\Gamma(t)$:
@@ C.1 Proof of Theorem 3
$\rho\chi-\frac{\rho\chi}{\Gamma_{t}}\Gamma(t)-\mathbb{E}^{\pi^{*}(t)}\left[H(q)\right]\left(\frac{\kappa^{\prime}(t)}{\chi}+\frac{\rho\Gamma_{t}^{\prime}}{\Gamma_{t}^{2}}\right)\Gamma(t)+$
$\left(\mathbb{E}^{\pi^{*}(\kappa(t))}\left[\widehat{\mu}(q)-\left(\frac{\kappa(t)}{\chi}-\frac{\rho}{\Gamma_{t}}\right)H(q)\right]\right)\gamma_{t}=0$
Note that the IC implies $\left(\mathbb{E}^{\pi^{*}(\kappa(t))}\left[\widehat{\mu}(q)-\frac{\kappa(t)}{\chi}H(q)\right]\right)=0$. Reorganizing terms, we get:
$\gamma(t)=\frac{\Gamma_{t}}{\rho\mathbb{E}^{\pi^{*}(\kappa(t))}[H(q)]}\left(\rho\chi\left(\frac{\Gamma(t)}{\Gamma_{t}}-1\right)+\mathbb{E}^{\pi^{*}(t)}\left[H(q)\right]\left(\frac{\kappa^{\prime}(t)}{\chi}+\frac{\rho\Gamma_{t}^{\prime}}{\Gamma_{t}^{2}}\right)\Gamma(t)\right).$
The ODE with initial conditions $\gamma(T)=0$, $\Gamma(T)=\Gamma_{T}$ satisfies the Lindelof condition and has a unique solution. Note that when $\Gamma(t)=\Gamma_{t}$, $\gamma(t)=\frac{\kappa^{\prime}(t)}{\rho\chi}\Gamma_{t}^{2}+\Gamma^{\prime}_{t}>0$, so $\Gamma(t)$ stays below $\Gamma_{t}$. When $\Gamma(t)=0$, $\gamma(t)<0$, so $\Gamma(t)$ stays above 00. Therefore, there exists $\underline{t}<T$ s.t. on $[\underline{t},T]$, $\Gamma(t)\geq 0$ and weakly increases.
By the construction of $\underline{t}$, for all $t_{0}\geq\underline{t}$, there exist multipliers $\lambda(t)$ and $\gamma(t)$ on $[t_{0},T]$ s.t. the Lagrangian $\mathcal{L}$ is maximized jointly by $(\lambda(t),\gamma(t))$ and $F^{*}$. ∎
C.2 Proof of Lemma 3
@@
By definition $\tau^{\prime}\leq\tau$. Since $\mathrm{Supp}(P)\subset D^{d}$ and $\mathrm{Supp(q_{\tau^{\prime}-})}\subset E^{A}$, $\mathrm{Supp}(q_{\tau^{\prime}})\subset\bar{Q}^{d}$. We prove by contradiction that $\tau^{\prime}=\tau$. Suppose $\tau^{\prime}<\tau$ on a positive measure, on which
$\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-(\tau-\tau^{\prime})\cdot\bar{\kappa}\big{|}\tau^{\prime}<\tau\right]=$ $\mathbb{E}^{P}\left[\mathbb{E}\left[\widehat{u}(q_{\tau})-(\tau-\tau^{\prime})\cdot\bar{\kappa}\big{|}\tau^{\prime}\right]\big{|}\tau^{\prime}<\tau\right]$
@@
The inequality is from the fact that $\tau^{\prime}<\tau\implies q_{\tau^{\prime}}\not\in E^{A}$. Therefore, $\tau^{\prime}$ strictly improves upon $\tau$ and $\tau$ is not incentive compatible. The contradiction implies that $\mathrm{Supp}(q_{\tau})\subset\bar{Q}^{d}$. ∎
C.3 Proof of Proposition 10
@@
As is discussed in Section 3, the agent-optimal policy can be solved by concavifying $\widehat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)$. Therefore, there exists a linear function $L(q)$ that is weakly higher than $\widehat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)$ and tangents it at two beliefs $q^{1}<\bar{q}_{0}<q^{2}$. WLOG, let $q^{1}$ and $q^{2}$ be the smaller and largest such beliefs, respectively. Since $\widehat{u}-\frac{\bar{\kappa}}{\chi}H$ is piece-wise strictly concave, the interval $[q^{1},q^{2}]$ is bounded away from the rest of $E^{A}$.
Proposition 3 implies that any admissible principal’s strategy has $\mathrm{Supp}(q_{\tau})\subset\bar{Q}^{0}$, which is the closure of $E^{A}$. Moreover, any continuous path that starts from $\bar{q}_{0}$ and ends outside of $[q^{1},q^{2}]$ leaves $E^{A}$; hence, it is not admissible. Therefore, $\mathrm{Supp}(q_{\tau})\subset[q^{1},q^{2}]$.
@@
Since $G$ is strictly convex, the relaxed problem is solved by $\pi^{*}$ with support $\left\{q^{1},q^{2}\right\}$ (such $\pi^{*}$ is unique). By Hébert and Woodford [2023], there exists a Gaussian process that implements $\pi^{*}$ and satisfies the information constraint. Note that this Gaussian process also implements the agent maximal continuation payoff (the upper concave hull of $\widehat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q)-H(q_{t}))$) for every interim belief; hence, it is incentive compatible. ∎

### v6

@@ A.7 Proof of ˜5
We slightly abuse notation and let $x_{1},x_{2}$ denote the random variables that determines the state realization. $\forall\pi$, let $\pi_{i}=\pi|_{X_{i}}$. Let $s_{1}$ be the random variable that is independent to $x_{2}$ and induces posterior belief $\pi_{1}$ for $x_{1}$. Let $s_{2}$ be the random variable that induces posterior belief $\pi$. Let $I$ denote mutual information. Then,
@@
In the derivation, $(q,\bar{q}_{0}|_{x_{2}})$ denotes the belief vector where the probability of state $x_{1}$ is given by $q$ and the probability of state $x_{2}$ is given by $\bar{q}_{0}$ restricted to $x_{2}$, respectively. Belief vector $(\bar{q}_{0}|_{X_{1}},q)$ is defined analogously. Therefore, let $\pi^{\prime}=\pi_{1}\otimes\pi_{2}$,
@@
$\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi_{1}}[\widehat{u}(q,\bar{q}_{0}|_{X_{2}})-\widehat{u}(\bar{q}_{0})]=\mathbb{E}^{\pi^{\prime}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})];$
@@
$\mathbb{E}^{\pi_{1}\otimes\pi_{2}}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)+\lambda G(q)\right]$
$=$ $\underbrace{\mathbb{E}^{\pi_{1}}\left[\widehat{u}(q,\bar{q}_{0}|_{X_{2}})-\frac{\kappa}{\chi}H(q,\bar{q}_{0}|_{X_{2}})\right]}_{\text{maximized by $\pi^{*}_{1}$ solving $V^{B}$}}+\mathbb{E}^{\pi_{2}}\left[-\frac{\kappa}{\chi}H(\bar{q}_{0}|_{X_{1}},q)+\lambda G(\bar{q}_{0}|_{X_{1}},q)\right].$
Appendix B Omitted Proofs for Section˜4
B.1 Proof of ˜6
@@
By Theorem 1, there is some $\pi\in\Pi^{*}$ that is the law of $q_{\tau}$ under the optimal policy. By the martingale property of the belief process, $q_{t}=E^{P}[q_{\tau}|\mathcal{F}_{t},t<\tau]$ for all $t\in[0,\tau)$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi$; hence, $\overline{V}_{R}(q_{t},\pi)$ is well defined.
For the purpose of contradiction, suppose that the inequality in Proposition 6 does not hold. Then, there exists $\epsilon>0$ and a positive probability event $\Omega^{\prime}\in\mathcal{F}_{t}|_{t<\tau}$ s.t. $\forall\omega\in\Omega^{\prime}$, $\overline{V}_{R}(q_{t}(\omega),\pi)<-\epsilon$. By the definition of $\overline{V}$,
@@
$\mathbb{E}\left[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\frac{\kappa}{\chi}(H(q_{\tau})-H(q_{t}))\Big|t<\tau,\omega\in\Omega^{\prime}\right]\leq-\epsilon.$
However, this violates the incentive compatibility condition of the agent. By deviating to stopping immediately at event $\Omega^{\prime}$, the agent’s expected utility improves by $\epsilon\cdot\mathrm{Prob}^{P}(\Omega^{\prime})$. Contradiction. ∎
B.2 Proof of Lemma 3
@@
By definition $\tau^{\prime}\leq\tau$. Let $\bar{E}^{A}$ denote the closure of $E^{A}$. Since $\mathrm{Supp(q_{\tau^{\prime}-})}\subset E^{A}$, $\mathrm{Supp}(q_{\tau^{\prime}})\subset\bar{E}^{A}$. We prove by contradiction that $\tau^{\prime}=\tau$. Suppose $\tau^{\prime}<\tau$ on a positive measure, on which
$\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-(\tau-\tau^{\prime})\cdot\kappa\big|\tau^{\prime}<\tau\right]=$ $\mathbb{E}^{P}\left[\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-(\tau-\tau^{\prime})\cdot\kappa\big|\tau^{\prime}\right]\big|\tau^{\prime}<\tau\right]$
@@
The inequality is from the fact that $\tau^{\prime}<\tau\implies q_{\tau^{\prime}}\not\in E^{A}$. Therefore, $\tau^{\prime}$ strictly improves upon $\tau$; hence, $\tau$ is not incentive compatible. The contradiction implies that $\mathrm{Supp}(q_{\tau})\subset\bar{E}^{A}$. ∎
B.3 Proof of Proposition 7
@@
As is discussed in Section 3, the agent-optimal policy can be solved by concavifying $\widehat{u}(q)-\frac{\kappa}{\chi}H(q)$. Therefore, there exists a linear function $L(q)$ that is weakly higher than $\widehat{u}(q)-\frac{\kappa}{\chi}H(q)$ and tangents it at two beliefs $q^{1}<\bar{q}_{0}<q^{2}$. WLOG, let $q^{1}$ and $q^{2}$ be the smaller and largest such beliefs, respectively. Since $\widehat{u}-\frac{\kappa}{\chi}H$ is piece-wise strictly concave, the interval $q^{1},q^{2}]$ is bounded away from the rest of $E^{A}$.
Proposition 3 implies that any admissible principal’s strategy has $\mathrm{Supp}(q_{\tau})\subset\bar{E}^{A}$. Moreover, any continuous path that starts from $\bar{q}_{0}$ and ends outside of $q^{1},q^{2}]$ leaves $E^{A}$; hence, it is not admissible. Therefore, $\mathrm{Supp}(q_{\tau})\subset[q^{1},q^{2}]$.
@@
Suppose $G$ is affine on $q^{1},q^{2}]$, then the principal is completely indifferent between any strategy; hence, the agent optimal strategy is optimal for the principal as well. If $G$ is not affine on $q^{1},q^{2}]$, the relaxed problem is solved by $\pi^{*}$ with support $\left\{q^{1},q^{2}\right\}$ (such $\pi^{*}$ is unique). By hebert2019rational, there exists a Gaussian process that implements $\pi^{*}$ and satisfies the information constraint. Note that this Gaussian process also implements the agent maximal continuation payoff (the upper concave hull of $\widehat{u}(q)-\frac{\kappa}{\chi}(H(q)-H(q_{t}))$) for every interim belief; hence, it is incentive compatible. ∎
B.4 Proof of ˜2
Proof.
Sufficiency: condition (ii) of Definition 3 is a direct implication of Propositions 1 and 2. Next, we verify condition (i). Suppose for the purpose of contradiction that condition (i) is violated, i.e. there exist some $t$, a positive probability Borel set $B\in\mathcal{F}_{t}$ and $\epsilon>0$ s.t. $\forall B^{\prime}\subset B$ and $B^{\prime}\in\mathcal{F}_{t}$,
$\mathbb{E}^{P}[G(q_{\tau})-G(q_{t})|\omega\in B^{\prime}]\leq\mathbb{E}^{P}[J(q_{t})|\omega\in B^{\prime}]-\epsilon;$
Let $\pi^{\prime}_{q}$ be the policy that implements $J(q)$ for each $q$. Let $\pi$ be the distribution of $q_{\tau}$. Then, the inequality above implies
$\mathbb{E}^{P}[\mathbb{E}^{\pi^{\prime}_{q_{t}(\omega)}}[G(q)|\omega\in B]P(B)+\mathbb{E}^{P}[G(q_{\tau})|\omega\not\in B](1-P(B))>\mathbb{E}^{\pi}[G(q)].$
Meanwhile, (4) implies that
$\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\kappa(\tau-t)|\omega\in B^{\prime}]=0.$
Since each $\pi_{q}$ is incentive compatible itself,
$\mathbb{E}^{P}[\mathbb{E}^{\pi^{\prime}_{q_{t}(\omega)}}[\widehat{u}(q)-\widehat{u}(q_{t}(\omega))-\frac{\kappa}{\chi}(H(q)-H(q_{t}(\omega))]|\omega\in B]$
$\geq$ $0=\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\frac{\kappa}{\chi}(H(q_{\tau})-H(q_{t}))|\omega\in B]$
[…]
The two inequalities above implies that the probability measure $\pi^{\prime\prime}=\mathbb{E}^{P}[\pi^{\prime}_{q_{t}(\omega)}|\omega\in B]P(B)+\mathbb{E}^{P}[\delta_{q_{\tau}}|\omega\not\in B](1-P(B))$ is feasible in (R) and strictly improves upon $\pi$, leading to a contradiction.
Necessity: Suppose for the purpose of contradiction that the equality doesn’t hold. Then, there exist some $t$, a positive probability Borel set $B\in\mathcal{F}_{t}$ and $\epsilon>0$ s.t. $\forall B^{\prime}\subset B$ and $B^{\prime}\in\mathcal{F}_{t}$,
$\mathbb{E}^{P}[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\kappa(\tau-t)|\omega\in B^{\prime}]\geq\epsilon$
$\implies$ $\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\frac{\kappa}{\chi}(H(q_{\tau})-H(q_{t}))|\omega\in B^{\prime}\right]\geq\epsilon.$
Let $\bar{I}=\max_{q\in\mathcal{P}(X)}(\sum q_{x}H(e_{x})-H(q))$. Let $\delta=\bar{I}/(\bar{I}+\epsilon)$. Let $\pi^{\prime}_{q}$ be the distribution of $q_{\tau}|_{q_{t}=t}$ for $q\in\mathrm{supp}q_{t}$. Let $\pi^{\prime\prime}_{q}=\delta\pi^{\prime}_{q}+(1-\delta)\sum q_{x}e_{x}$. Then, by construction, $\forall B^{\prime}\subset B$,
$\mathbb{E}^{P}\left[\mathbb{E}^{\pi^{\prime\prime}_{q_{t}(\omega)}}\left[\widehat{u}(q)-\widehat{u}(q_{t})-\frac{\kappa}{\chi}(H(q)-H(q_{t}))\right]\Big|\omega\in B^{\prime}\right]\geq 0.$
Since $\pi^{\prime\prime}_{q}$ leaves the agent with non-negative utility, it is feasible in (R); hence, $\mathbb{E}^{\pi^{\prime\prime}_{q_{t}(\omega)}}[G(q)-G(q_{t}(\omega))]\leq J(q_{t}(\omega))$. By assumption, $\mathrm{supp}(q_{\tau})\subset\mathcal{P}(X)^{+}$; hence, $\mathrm{supp}(q_{t})\subset\mathcal{P}(X)^{+}$ due to the martingale property. Therefore, $\forall B^{\prime}\subset B$, $\mathbb{E}^{P}[G(q_{\tau})|\omega\in B^{\prime}]<\mathbb{E}^{P}[\sum q_{x}(\omega)G(e_{x})|\omega\in B^{\prime}]$. This implies,
$\mathbb{E}^{P}[G(q_{\tau})-G(q_{t}(\omega))|\omega\in B^{\prime}]<\mathbb{E}^{P}\left[\mathbb{E}^{\pi^{\prime\prime}_{q_{t}(\omega)}}\left[G(q)-G(q_{t}(\omega))\right]\big|\omega\in B^{\prime}\right]\leq\mathbb{E}^{P}[J(q_{t})|\omega\in B^{\prime}].$
This violates condition (i), which states
$\mathrm{Prob}^{P}\Big(\mathbb{E}^{P}[G(q_{\tau})-G(q_{t})-J(q_{t})|\mathcal{F}_{t}]\geq 0\Big)=1.$
B.5 Proof of ˜8
Proof.
By Theorem 1, there is some $\pi\in\Pi^{*}$ that is the law of $q_{\tau}$ under the optimal policies. By the martingale property of the belief process, $q_{t}=E^{P}[q_{\tau}|\mathcal{F}_{t},t<\tau]$ for all $t\in[0,\tau)$, and thus $q_{t}$ must lie in the convex hull of the support of $\pi$.
For the sake of contradiction, suppose Proposition 8 does not hold. Then, there exists positive probability event $B^{\prime}\in\mathcal{F}_{t}|_{t<\tau}$ and $\epsilon>0$ such that $\forall\omega\in B^{\prime}$,
$\inf_{\pi\in\Pi^{*}}\underline{V}_{R}(q_{t}(\omega),\pi)\geq\epsilon.$
Then, $\mathbb{E}^{P}\left[\widehat{u}(q_{\tau})-\widehat{u}(q_{t})-\frac{\kappa}{\chi}(H(q_{\tau})-H(q_{t}))\Big|\omega\in B^{\prime}\right]\geq\epsilon,$ i.e., the agent obtains positive interim surplus on positive probability event $B^{\prime}$, which violates ˜2. ∎